OpenStudy (anonymous):

whats x? 2x+3(x-4)=18

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Distribute the 3 into the parenthesis 2x + 3x - 12 = 18 Combine like terms. 2x +3x - 12 = 18 +12 +12 5x = 30 /5 /5 x=6

OpenStudy (anonymous):

x=6